# Capturing Temporal Dynamics in Large-Scale Canopy Tree Height Estimation

### v1

@@ Abstract
With the rise in global greenhouse gas emissions, accurate large-scale tree canopy height maps are essential for understanding forest structure, estimating above-ground biomass, and monitoring ecological disruptions. To this end, we present a novel approach to generate large-scale, high-resolution canopy height maps over time. Our model accurately predicts canopy height over multiple years given Sentinel 2 time series satellite data. Using GEDI LiDAR data as the ground truth for training the model, we present the first 10 m resolution temporal canopy height map of the European continent for the period 2019–2022. As part of this product, we also offer a detailed canopy height map for 2020, providing more precise estimates than previous studies. Our pipeline and the resulting temporal height map are publicly available, enabling comprehensive large-scale monitoring of forests and, hence, facilitating future research and ecological analyses. For an interactive viewer, see https://europetreemap.projects.earthengine.app/view/temporalcanopyheight.
@@ 1 Introduction
Recent studies have advanced canopy height prediction using classical machine learning methods (Potapov et al., 2021; Kacic et al., 2023) and deep learning approaches (Schwartz et al., 2024; Fayad et al., 2023; Lang et al., 2023; Pauls et al., 2024; Liu et al., 2023; Tolan et al., 2024), with the satellite-dependent resolution varying between 30 m (Landsat), 10 m (Sentinel-1, -2), 3 m (PlanetLabs) and 60 cm (Maxar) with only the Sentinel and Landsat program being openly available. However, almost all studies focus on predicting canopy height for a single year, despite that being insufficient (Bevacqua et al., 2024). Temporal dynamics of forests are essential for identifying carbon sources and sinks and assessing forest responses to natural or human-induced stressors such as drought, diseases, or environmental changes, and potentially predicting future forest developments. Although some methods have explored temporal tree height mapping222We define temporal maps as those which provide canopy heights over multiple years, independent of the generation process. at local scales (Kacic et al., 2023) and Turubanova et al. (2023) produced a map for Europe, no approach has yet tackled this challenge at a resolution of 10 m and greater scale combined, which is vital for assessing forest structural details.
In this work, we introduce an approach for generating large-scale, temporal tree height maps from satellite imagery. Specifically, we use a custom deep learning approach that predicts tree canopy height from Sentinel image data, using the sparsely-distributed space-borne GEDI LiDAR data as a ground truth source for tree height.
@@
We demonstrate that using a full 12-month time series of Sentinel-2 imagery, rather than using a single aggregated composite, yields substantial performance gains by allowing the model to capture seasonal patterns and leverage geolocation shifts.
The entire pipeline and model weights are publicly available on GitHub333https://github.com/AI4Forest/Europe-Temporal-Canopy-Height. Further, the resulting canopy height maps are accessible through Google’s Earth Engine (Gorelick et al., 2017), ensuring reproducibility and facilitating research on large-scale forest monitoring and above-ground biomass estimation. Our framework highlights the importance of using multi-temporal satellite data for forest monitoring, enabling precise, large-scale assessment of forest structure and dynamics.
@@ 2.1 Satellite Imagery
To further address these issues, temporal composites are used to aggregate images over time, employing techniques like per-pixel median calculation (Pauls et al., 2024; Schwartz et al., 2024; Fayad et al., 2023) or the Best-Available-Pixel approach (Senf & Seidl, 2021). These methods mitigate adverse weather and atmospheric effects, ensuring more consistent and reliable images for tree canopy height analysis. However, temporal aggregation often results in significant information loss, discarding valuable seasonal variations such as leaf-on vs. leaf-off seasons. This hinders capturing crucial temporal dynamics, but has thus far only been tackled for crop yield prediction and classification (Fan et al., 2021; Johnson & Mueller, 2021).
Additionally, aggregation removes the spatial shifts caused by geolocation inaccuracies, thus losing the opportunity to leverage this variability for more accurate predictions. Precisely, Sentinel-2 imagery exhibits a mean geolocation offset of about 4 m (Yan et al., 2018), causing each pixel’s reflectance to be influenced by its surroundings. Leveraging this offset across multiple images can enhance the model’s ability to detect structural details, such as distinguishing a large tree’s borders within a forest of smaller trees. This results in more precise height predictions compared to using aggregated images. Wolters et al. (2023) demonstrated that using multiple Sentinel-2 images outperforms single-image methods by leveraging temporal information and geolocation shifts. Notably, their model remained effective even with cloudy inputs, suggesting it can focus on clear portions of the data, enhancing robustness. Our work further builds upon this approach.
@@ 2.2 Tree Canopy Height Estimation
A key challenge in training models for temporal canopy height estimation is obtaining accurate ground-truth data. This data can come from national forest inventories or LiDAR measurements, either airborne or spaceborne. Airborne LiDAR provides high-resolution data but is limited to local areas. In contrast, NASA’s GEDI mission offers broader geographic and temporal coverage with spaceborne LiDAR, though at a coarser spatial resolution and with occasional geolocation inaccuracies. GEDI measures canopy heights using laser shots with a 25 m footprint, but only about 4% of Earth’s surface is covered during the satellite’s operational period, and spatial alignment across different years is virtually non-existent. This inconsistency complicates the use of GEDI as a reliable temporal training dataset.
In consequence, most studies focus on estimating height for a single year, with few addressing multi-year time series (Dixon et al., 2025; Kacic et al., 2023; Turubanova et al., 2023). Mapping efforts over multiple years or seasons face significant computational and storage challenges and can result in fluctuating height estimates due to varying satellite conditions and model uncertainties. A common approach is to apply a single-year model to data from other years, but this fails to capture temporal patterns. Consequently, post-processing techniques like moving averages, trend detection, and cut-detection are used to smooth out year-to-year variations. Developing a robust, large-scale framework for temporal tree canopy height mapping remains a key research challenge in forest monitoring.
@@ 3 Approach
Our methodology integrates multi-source satellite imagery and GEDI LiDAR data to produce high-resolution, temporal canopy height maps for Europe. This section details the data sources, preprocessing steps, and the design of our deep learning framework.
@@ 3.1 Data
To construct our dataset, we integrate information from three key sources: Sentinel-1, Sentinel-2, and GEDI. Each dataset undergoes rigorous preprocessing to ensure compatibility and optimal quality. To keep the best projection accuracy, we use a tiling system following the Universal Transverse Mercator (UTM)444The UTM coordinate system divides the Earth into 120 zones, each of longitude wide, using transverse Mercator projections to minimize distortion. Each zone is further divided into tiles following the Sentinel-2 tiling system for better data handling. system and process the data in these tiles.
@@
Sentinel-2 multispectral imagery is a key part of our approach. We use the Level-2A surface reflectance product (BOA) available via the Copernicus AWS, selecting a single best image each month based on minimal cloud cover, thereby ensuring minimal contamination and consistent data quality. This yields a temporal sequence of 12 monthly images per tile and year (one per month), thereby preserving seasonal patterns important for vegetation monitoring.
We include all Sentinel-2 bands except B10 (cirrus). To prepare the training data, each band’s reflectance values are normalized by a fixed factor (cf. Table 1), mapping values to . Rather than employing conventional min-max or zero-mean normalization techniques, which are substantially affected by the presence of clouds and cirrus, we identified value ranges for each band that contain “valuable” information, ensuring data quality is maintained for analysis. By retaining monthly variability instead of aggregating temporal data, our approach captures seasonal dynamics, such as transitions between leaf-on and leaf-off states, which are crucial for vegetation monitoring.
| Bands | Normalization Divisor |
@@ 3.2 Model Architecture
We adopt a three-dimensional (3D) U-Net inspired by Çiçek et al. (2016), originally proposed for volumetric segmentation tasks, and modify it to output a single-channel tree canopy height for each input pixel. Unlike prior approaches relying on static images or single composites (Schwartz et al., 2024; Pauls et al., 2024), our model processes a stack of 12 monthly Sentinel-2 images concatenated with an aggregated Sentinel-1 composite. This approach exploits temporal patterns (e.g., seasonal vegetation changes), and leverages geolocation offsets in Sentinel-2 imagery and enhances spatial resolution by incorporating information from neighboring pixels.
@@ 3.3 Model Training
We train our models using Adam (Kingma, 2014) with an initial learning rate of 0.001, weight decay of 0.01, and gradient clipping at 1.0. We follow best practices and use a linear learning rate scheduler with a 10% warmup, training for iterations with a batch size of 16 (corresponding to 8 epochs).
@@ 4 Results
This section presents the results of our temporal tree canopy height estimation model, including quantitative and qualitative evaluations. We focus on the accuracy of our predicted tree canopy height maps, comparisons with existing models, and generalization across diverse european forest regions.
@@ 4.2 Quantitative Evaluation
We begin by comparing our different model configurations for the reference year 2020 across different metrics: Mean Absolute Error (MAE), Mean Squared Error (MSE) and Root Mean Squared Error (RMSE). As the resulting map is targeted towards forest assessment, we compare it only against labels exceeding , to exclude labels over acres and grassland. Table 2 excludes 2D-Composite-Year, 2D-Stack-Year, and 3D-Stack-Year to avoid redundancy, as 2D-Composite-Year is the same as 2D-Composite-2020 for the 2020 comparison.
| MAE [m] | MSE [m2] | RMSE [m] | |
| 2D-Composite-2020 | 5.66 | 85.93 | 7.60 |
| 2D-Composite-MultiYear | 5.43 | 82.70 | 7.37 |
| 2D-Stack-2020 | 5.51 | 81.31 | 7.37 |
| 2D-Stack-MultiYear | 5.17 | 78.07 | 7.08 |
| 3D-Stack-2020 | 5.48 | 80.32 | 7.29 |
| 3D-Stack-MultiYear | 5.05 | 77.40 | 7.01 |
@@ 4.3 Method Comparison
| Source | MAE [m] | MSE [m2] | RMSE [m] | |
| Tolan et al. (2024) | Maxar | 11.25 | 212.14 | 13.25 |
| Liu et al. (2023) | Planet | 8.17 | 138.25 | 10.36 |
| Lang et al. (2023) | S2 | 5.74 | 84.68 | 7.57 |
| Pauls et al. (2024) | S1/2 | 5.46 | 83.14 | 7.40 |
| Turubanova et al. | Landsat | 12.39 | 252.57 | 14.14 |
| Ours | S1/2 | 4.76 | 74.28 | 6.75 |
Table 4 reports MAE, MSE, and RMSE for five maps: a global map at 1 m resolution (Tolan et al., 2024), a European map at 3 m resolution (Liu et al., 2023)555Liu et al. (2023)’s 2019 map is included for completeness., global maps at 10 m resolution (Lang et al., 2023; Pauls et al., 2024), and a European map at 30 m resolution (Turubanova et al., 2023).
Despite the coarser resolution of Sentinel-1/2 (10 m) compared to Planet (3 m) or Maxar (60 cm), models utilizing Sentinel data achieve higher accuracy, likely due to the availability of near-infrared and shortwave-infrared bands. As shown in Table 4, our model, 3D-Stack-MultiYear-L, achieves the best performance across all metrics, with an MAE of 4.76 m—representing a 13% improvement over the next-best model (Pauls et al., 2024). Similar improvements are observed for MSE and RMSE.
@@
Figure 5 shows a scatterplot for each model along with its value, where (coefficient of determination) quantifies the proportion of variance in the tree canopy height labels explained by the model predictions. An value of 1 indicates a perfect fit, while means no variance is explained. Here, is calculated using all labels and takes only labels exceeding .
The plots reveal that Tolan et al. (2024) and Turubanova et al. (2023) struggle with taller trees, as neither predict heights beyond . Lang et al. (2023), Liu et al. (2023), and Pauls et al. (2024) tend to saturate at . In contrast, our model exhibits the narrowest point cloud, with the highest density areas aligning closely with the perfect fit line, and achieves strong performance up to –. Consequently, we improve from to , and for labels over (), from to .
@@
While quantitative evaluation is essential, visual quality is equally important for tree canopy height prediction. Figure 4 visually compares Liu et al. (2023), Tolan et al. (2024), Lang et al. (2023), Pauls et al. (2024), Turubanova et al. (2023) and our model. The first column shows a high-resolution satellite image from Google Maps (not necessarily from 2020) whereas all other columns show the height between (black) and (light yellow). Despite its relatively modest metrics, the high resolution of Tolan et al. (2024)’s map makes it effective for canopy detection. Liu et al. (2023) identifies trees with higher resolution and does not suffer as much from saturation effects as Tolan et al. (2024) and Turubanova et al. (2023). While our model lacks the high resolution of Liu et al. (2023) and Tolan et al. (2024), it surpasses Pauls et al. (2024), Lang et al. (2023), and Turubanova et al. (2023). Unlike other maps that assign similar heights to all forest patches, our map is the only one that differentiates between various forest heights in smaller patches (lower part of the first image row).
@@
As our model is applied to 2019-2022 data, we can observe temporal dynamics such as deforestation, identifying where trees have been cut down. Figure 6 shows observable deforestation in two areas. However, mapping minor tree growth is challenging due to the short time frame and the slow natural growth rate of trees. In contrast, in forest plantation areas where fast-growing tree species are actively managed, significant growth can be observed easier because the acceleration in growth outweighs the uncertainty in predictions.
@@ 5 Conclusion
We presented a novel approach for generating high-resolution, large-scale temporal tree canopy height maps using a 3D U-Net model. By utilizing a full 12-month Sentinel-2 imagery time series, our approach avoids the information loss of median composites and captures essential seasonal variations and geolocation shifts. Trained on GEDI LiDAR data, our model produces a highly accurate 10 m resolution tree canopy height map of Europe for the years 2019 to 2022. Our model outperforms existing state-of-the-art models, achieving significantly lower errors. Notably, it drastically improves accuracy for tall trees, essential for precise biomass estimation and carbon stock assessments. Our temporal maps capture forest dynamics, such as deforestation and growth patterns, offering valuable insights for forest monitoring and ecological analyses. The publicly released pipeline and tree canopy height maps will support further research and inform decision-making in forest management and conservation. Our work underscores the importance of multi-temporal satellite data in enhancing the accuracy and reliability of large-scale forest assessments.

### v2

@@ Abstract
With the rise in global greenhouse gas emissions, accurate large-scale tree canopy height maps are essential for understanding forest structure, estimating above-ground biomass, and monitoring ecological disruptions. To this end, we present a novel approach to generate large-scale, high-resolution canopy height maps over time. Our model accurately predicts canopy height over multiple years given Sentinel-1 composite and Sentinel 2 time series satellite data. Using GEDI LiDAR data as the ground truth for training the model, we present the first 10 m resolution temporal canopy height map of the European continent for the period 2019–2022. As part of this product, we also offer a detailed canopy height map for 2020, providing more precise estimates than previous studies. Our pipeline and the resulting temporal height map are publicly available, enabling comprehensive large-scale monitoring of forests and, hence, facilitating future research and ecological analyses.
@@ 1 Introduction
Recent studies have advanced canopy height prediction using classical machine learning methods (Potapov et al., 2021; Kacic et al., 2023; Loo & Wang, 2024) and deep learning approaches (Schwartz et al., 2024; Fayad et al., 2023; Lang et al., 2023; Pauls et al., 2024; Liu et al., 2023; Tolan et al., 2024; Fayad et al., 2025), with the satellite-dependent resolution varying between 30 m (Landsat), 10 m (Sentinel-1, -2), 3 m (PlanetLabs) and 60 cm (Maxar) with only the Sentinel and Landsat program being openly available. However, almost all studies focus on predicting canopy height for a single year, despite that being insufficient (Bevacqua et al., 2024). Temporal dynamics of forests are essential for identifying carbon sources and sinks and assessing forest responses to natural or human-induced stressors such as drought, diseases, or environmental changes. Although some methods have explored temporal tree height mapping222We define temporal maps as those which provide canopy heights over multiple years, independent of the generation process. at local scales (Kacic et al., 2023) and Turubanova et al. (2023) produced a map for Europe, no approach has yet tackled this challenge at a resolution of 10 m and greater scale combined, which is vital for assessing forest details.
In this work, we introduce an approach for generating large-scale, temporal tree height maps. Specifically, we use a custom deep learning approach that predicts tree canopy height from Sentinel image data, using the sparsely-distributed space-borne GEDI LiDAR data as a ground truth.
@@
We demonstrate that using a 12-month time series of Sentinel-2 imagery, rather than using a single aggregated composite, yields substantial performance gains.
The entire pipeline and model weights are publicly released on GitHub333https://github.com/AI4Forest/Europe-Temporal-Canopy-Height and the resulting tree canopy height maps are accessible through Google’s Earth Engine444https://europetreemap.projects.earthengine.app/view/europeheight (Gorelick et al., 2017), ensuring reproducibility and facilitating research on large-scale forest monitoring, forest structure analysis and above-ground biomass estimation.
@@ 2.1 Satellite Imagery
To further address these issues, temporal composites are used to aggregate images over time, employing techniques like per-pixel median calculation (Pauls et al., 2024; Schwartz et al., 2024; Fayad et al., 2023) or the Best-Available-Pixel approach (Senf & Seidl, 2021). These methods mitigate adverse weather and atmospheric effects, ensuring more consistent and reliable images for tree canopy height analysis. However, temporal aggregation often results in significant information loss, discarding valuable seasonal variations such as leaf-on vs. leaf-off seasons. This hinders capturing crucial temporal dynamics, but has thus far only been tackled for crop yield prediction and classification (Fan et al., 2021; Johnson & Mueller, 2021) and general satellite image processing (Tarasiou et al., 2023).
Additionally, aggregation removes the spatial shifts caused by geolocation inaccuracies, thus losing the opportunity to leverage this variability for more accurate predictions. Precisely, Sentinel-2 imagery exhibits a mean geolocation offset of about 4 m (Yan et al., 2018), causing each pixel’s reflectance to be influenced by its surroundings. Leveraging this offset across multiple images can enhance the model’s ability to detect structural details, such as distinguishing a large tree’s borders within a forest of smaller trees. This results in more precise height predictions compared to using aggregated images. Wolters et al. (2023) demonstrated that using multiple Sentinel-2 images outperforms single-image methods by leveraging temporal information and geolocation shifts. Our work further builds upon this approach.
@@ 2.2 Tree Canopy Height Estimation
A key challenge in training models for temporal canopy height estimation is obtaining accurate ground-truth data. This data can come from national forest inventories or LiDAR measurements, either airborne or spaceborne. Airborne LiDAR provides high-resolution data but is limited to local areas. In contrast, NASA’s GEDI mission offers broader geographic and temporal coverage with spaceborne LiDAR, though at a coarser spatial resolution and with occasional geolocation inaccuracies. GEDI measures canopy heights555Although GEDI not only measures trees, this can be neglected here, as forest masks are used for filtering before further analysis. using laser shots with a 25 m footprint, but only about 4% of Earth’s surface is covered during the satellite’s operational period, and spatial alignment across different years is virtually non-existent. This inconsistency complicates the use of GEDI as a reliable temporal training dataset.
In consequence, most studies focus on estimating height for a single year, with few addressing multi-year time series (Dixon et al., 2025; Kacic et al., 2023; Turubanova et al., 2023). Mapping efforts over multiple years or seasons face significant computational and storage challenges and can result in fluctuating height estimates due to varying satellite conditions and model uncertainties. A common approach is to just apply a single-year model to data from other years, but this fails to capture temporal patterns. Consequently, post-processing techniques like moving averages, trend detection, and cut-detection are used to smooth out year-to-year variations. Developing a robust, large-scale framework for temporal tree canopy height mapping remains one of the key research challenges in forest monitoring.
@@ 3 Approach
Our methodology integrates multi-source satellite imagery and GEDI LiDAR data to produce high-resolution, temporal canopy height maps for Europe. Next, we detail the data sources, preprocessing steps, and the design choices made.
@@ 3.1 Data
To construct our dataset, we integrate information from three key sources: Sentinel-1, Sentinel-2, and GEDI. Each dataset undergoes rigorous preprocessing to ensure compatibility and optimal quality. To keep the best projection accuracy, we use a tiling system following the Universal Transverse Mercator (UTM)666The UTM coordinate system divides the Earth into 120 zones, each of longitude wide, using transverse Mercator projections to minimize distortion. Each zone is further divided into tiles following the Sentinel-2 tiling system for better data handling. system and process the data in these tiles.
@@
Sentinel-2 multispectral imagery is a key part of our approach. We use the Level-2A surface reflectance product (BOA) available via the Copernicus AWS, selecting a single best image each month based on minimal cloud cover, thereby ensuring minimal contamination and consistent data quality. This yields a temporal sequence of 12 monthly images per tile and year (one per month), thereby preserving seasonal patterns important for vegetation monitoring. We include all Sentinel-2 bands except B10 (cirrus). Each band’s reflectance values are normalized by a fixed factor (cf. Table 1), mapping values to . Rather than employing conventional min-max or zero-mean normalization techniques, which are substantially affected by the presence of clouds and cirrus, we identified value ranges for each band that contain “valuable” information, ensuring data quality is maintained for analysis. By retaining monthly variability instead of aggregating temporal data, our approach captures seasonal dynamics, such as transitions between leaf-on and leaf-off states, which are crucial for vegetation monitoring.
| Bands | Divisor |
@@ 3.2 Model Architecture
We adopt a three-dimensional (3D) U-Net inspired by Çiçek et al. (2016), originally proposed for volumetric segmentation tasks, and modify it to output a single-channel tree canopy height for each input pixel. Unlike prior approaches relying on static images or single composites (Schwartz et al., 2024; Pauls et al., 2024), our model processes a stack of 12 monthly Sentinel-2 images concatenated with an aggregated Sentinel-1 composite. This approach allows the model to use potential seasonal vegetation changes, and possibly exploit geolocation offset in Sentinel-2 imagery by incorporating information from neighboring pixels.
@@ 3.3 Model Training
We use Adam (Kingma, 2014) with an initial learning rate of 0.001, weight decay of 0.01, and gradient clipping at 1.0. We follow best practices (Li et al., 2020; Zimmer et al., 2023) and use a linear learning rate scheduler with a 10% warmup, training for iterations with a batch size of 16 (corresponding to 8 epochs).
@@ 4 Results
This section presents the results of our temporal tree canopy height model. We focus on the accuracy of our predicted tree canopy height maps, comparisons with existing models, and generalization across european forest regions.
@@ 4.2 Quantitative Evaluation
We begin by comparing our different model configurations for the reference year 2020 across two different metrics: Mean Absolute Error (MAE) and Mean Squared Error (MSE). As the resulting map is targeted towards forest assessment, we compare it only against labels exceeding , to exclude labels over acres and grassland. Table 2 excludes 2D-Composite-Year, 2D-Stack-Year, and 3D-Stack-Year to avoid redundancy, as 2D-Composite-Year is the same as 2D-Composite-2020 for the 2020 comparison.
| MAE [m] | MSE [m2] | |
| 2D-Composite-2020 | 5.66 | 85.93 |
| 2D-Composite-MultiYear | 5.43 | 82.70 |
| 2D-Stack-2020 | 5.51 | 81.31 |
| 2D-Stack-MultiYear | 5.17 | 78.07 |
| 3D-Stack-2020 | 5.48 | 80.32 |
| 3D-Stack-MultiYear | 5.05 | 77.40 |
@@ 4.3 Method Comparison
Table 4 reports MAE, MSE, and coefficient of determination for five maps: a global map at 1 m resolution (Tolan et al., 2024), a European map at 3 m resolution (Liu et al., 2023)777Liu et al. (2023)’s 2019 map is included for completeness., global maps at 10 m resolution (Lang et al., 2023; Pauls et al., 2024), and a European map at 30 m resolution (Turubanova et al., 2023).
Despite the coarser resolution of Sentinel-1/2 (10 m) compared to Planet (3 m) or Maxar (60 cm), models utilizing Sentinel data achieve higher accuracy, likely due to the availability of near-infrared and shortwave-infrared bands. As shown in Table 4, our model, 3D-Stack-MultiYear-L, achieves the best performance across all metrics, with an MAE of 4.76 m—representing a 13% improvement over the next-best model (Pauls et al., 2024). Similar improvements are observed for MSE and .
@@
Figure 5 shows a scatterplot for each model along with its value, where (coefficient of determination) quantifies the proportion of variance in the tree canopy height labels explained by the model predictions (a corresponds to a perfect fit, a means that no variance is explained). Here, is calculated using all labels and takes only labels exceeding . The plots reveal that Tolan et al. (2024) and Turubanova et al. (2023) struggle with taller trees, as neither predict heights beyond . Lang et al. (2023), Liu et al. (2023), and Pauls et al. (2024) tend to saturate at . In contrast, our model exhibits the narrowest point cloud, with the highest density areas aligning closely with the perfect fit line, and achieves strong performance up to –. Consequently, we improve from to , and for labels over (), from to .
@@
While quantitative evaluation is essential, visual quality is equally important for tree canopy height prediction. Figure 4 visually compares Liu et al. (2023), Tolan et al. (2024), Lang et al. (2023), Pauls et al. (2024), Turubanova et al. (2023) and our model. The first column shows a high-resolution satellite image from Google Maps (not necessarily from 2020) whereas all other columns show the height between (black) and (light yellow). Despite its relatively modest metrics, the high resolution of Tolan et al. (2024)’s map makes it effective for canopy detection. Liu et al. (2023) identifies trees with higher resolution and does not suffer as much from saturation effects as Tolan et al. (2024) and Turubanova et al. (2023). While our model lacks the high resolution of Liu et al. (2023) and Tolan et al. (2024), it surpasses Pauls et al. (2024), Lang et al. (2023), and Turubanova et al. (2023). Unlike other maps that assign similar heights to all forest patches, our map is the only one differentiating various forest heights in smaller patches (bottom of first row).
@@
As our model is applied to 2019-2022, we can observe temporal dynamics such as deforestation, identifying where trees have been cut down. Figure 6 shows observable deforestation in two areas. However, mapping minor tree growth is challenging due to the short time frame and slow natural growth rate of trees. In contrast, in forest plantations where fast-growing tree species are actively managed, significant growth can be observed more easily because the acceleration in growth outweighs the uncertainty in our predictions.
@@ 5 Conclusion
We presented a novel approach for generating high-resolution, large-scale temporal tree canopy height maps using a 3D U-Net model. By utilizing a full 12-month Sentinel-2 imagery time series, our approach avoids the information loss of median composites and captures essential seasonal variations and geolocation shifts. Trained on GEDI LiDAR data, our model produces a highly accurate 10 m resolution tree canopy height map of Europe for the years 2019 to 2022. Our model outperforms existing state-of-the-art models, achieving significantly lower errors. Notably, it drastically improves accuracy for tall trees, essential for precise biomass estimation and carbon stock assessments. Our temporal maps capture forest dynamics, such as deforestation and growth patterns, offering valuable insights for forest monitoring and ecological analyses. We believe that the publicly released pipeline and tree canopy height maps will support further research and inform decision-making in forest management and conservation.